A spinner is divided into 8 equal sections numbered 1-8. what is the probability of spinning an even number and then a number greater than 5?
1
Expert's answer
2010-12-05T05:54:08-0500
The probability spining of even number is 4/8 = 1/2
The probability spinning of number greater than 5 is 3/8
The probability of these two events will come together is 1/2*3/8 = 3/16